# Hamilton cycles in random digraphs with minimum degree at least one

### v1

@@ 1 Introduction
$\lim_{n\to\infty}\mathbb{P}(D_{n,m}^{(\delta\geq 1)}\text{ is Hamiltonian})=\begin{cases}0&c_{n}\to-\infty.\\ e^{-e^{-c}/4}&c_{n}\to c.\\ 1&c_{n}\to\infty.\end{cases}$ (2)
@@ 1.1 Previous work
Previous work in this area has focused on $G_{n,m}^{(\delta\geq k)}$, where we condition on $G_{n,m}$ having minimum degree at least $k$. This work was inspired by the result of Komlós and Szemerédi [14]:
@@ 2 Proof of Theorem 1
We will base our proof on a 3-phase method used in various forms in [6], [8], [10], [12]. We first need to describe a useable model for $D_{n,m}^{(\delta\geq 1)}$. For a sequence ${\bf x}=(x_{1},x_{2},\ldots,x_{2m})\in[n]^{2m}$ we let $D_{\bf x}$ be the multi-digraph with vertex set $[n]$ and edge set $E_{\bf x}=\left\{(x_{2j-1},x_{2j}):j=1,2,\ldots,m\right\}$. Let $d^{+}_{{\bf x}}(i)=|\left\{j:x_{2j-1}=i\right\}|$ be the out-degree of $i\in[n]$ in $D_{\bf x}$ and similarly let $d^{-}_{{\bf x}}(i)=|\left\{j:x_{2j}=i\right\}|$ be the in-degree of $i\in[n]$ in $D_{\bf x}$. Then $\Omega_{1}=\left\{{\bf x}\in[n]^{2m}:d^{+}_{{\bf x}}(i),d^{-}_{{\bf x}}(i)\geq 1\right\}$ for all $i\in[n]$.
@@
For ${\bf x}\in\Omega$, let $L_{\bf x}=\left\{j\in[m]:x_{2j-1}=x_{2j}\right\}$ be the set of loops in $D_{\bf x}$, let $M_{\bf x}=\{j\in[m]:\exists j^{\prime}\neq j\text{ s.t. }(x_{2j^{\prime}-1},x_{2j^{\prime}})=$ $(x_{2j-1},x_{2j})\}$ define the set of multiple edges in $D_{\bf x}$; and let $\Delta_{\bf x}=\max_{j}(d_{{\bf x}}^{-}(j)+d_{{\bf x}}^{+}(j))$. (We show later, see Lemma 12 below, that if $m\leq n\log n$ then w.h.p. there are no multiple loops, no triple edges, no loops incident to double edges.)
@@
Consequently, choosing ${\bf x}$ (near) uniformly from $\Omega_{1}^{*}$ and taking $D_{\bf x}$ is equivalent to sampling (near) uniformly from ${\mathcal{D}}_{n,m}^{(\delta\geq 1)}$. By uniform sampling from $\Omega_{1}$ and then using switchings to remove any loops or parallel edges and we hence obtain a space $\Omega_{1}^{(0,0)}$ which is almost all of $\Omega_{1}^{*}$ and whose entries are (near) uniformly distributed $\Omega_{1}^{(0,0)}$.
We can generate a random member of $\Omega_{1}$ as follows: Let $Z=Z(z)$ be a truncated Poisson random variable with parameter $z$, where
@@
Let $Z_{i}^{\prime},Z_{i}^{\prime\prime}\,(j\in[n])$ be independent copies of a truncated Poisson random variable $Z$. Let ${\bf x}$ be chosen randomly from $\Omega_{1}$. Then $\{d_{\bf x}^{+}(i)\}_{j\in[n]}$ is distributed as $\{Z_{i}^{\prime}\}_{i\in[n]}$ conditional on $\sum_{j\in[n]}Z_{i}^{\prime}=m$ and similarly for $\{d_{\bf x}^{-}(i)\}_{i\in[n]}$ with respect to $Z_{i}^{\prime\prime}$.
@@
Given ${\bf x}\in\Omega_{1}^{(k,0)}$ with $k>0$, we will define an $L$-switch that removes a loop $(x,x)$ by choosing an edge $(a,b)$ and replacing the pair by $(x,a),(x,b)$. More formally, suppose that $i\in L_{\bf x}$. Let $j\notin L_{\bf x}$. Then an $L$-switch produces ${\bf x}^{\prime}$ where we make the replacements $x^{\prime}_{2i-1}\leftarrow x_{2j-1},x^{\prime}_{2j-1}\leftarrow x_{2i-1}$. After $k$ $P$-switches we will have produced a member of $\Omega_{1}^{(0,0)}$. (The above can make new parallel edges, but they are few in number and we show in Lemma 5 that we can exclude switches that create new parallel edges.)
These switches do not preserve the precise degree sequence $d_{{\bf x}}^{+}(i),d_{{\bf x}}^{-}(i),i\in[n]$ for ${\bf x}\in\Omega_{1}$, although they preserve the total degrees $d_{{\bf x}}^{+}(i)+d_{{\bf x}}^{-}(i),i\in[n]$ for ${\bf x}\in\Omega_{1}$. However, they do preserve $d_{{\bf x}}^{+}(i)\geq 1,d_{{\bf x}}^{-}(i)\geq 1$ w.h.p. for $i\in[n]$.
@@
With a little bit of effort, we can use the analysis for Claim 3 to obtain an estimate of ${\mathcal{D}}_{n,m}^{(\delta\geq 1)}$, see Appendix C.
@@ 2.1 Phase 0
Our input will be a sequence y constructed as in Claim 3, starting with a random ${\bf x}\in\Omega_{1}$. We will discuss Hamiltonicity in the context of $D_{\bf y}$. This phase partitions $E_{\bf y}$ into three subsets $E_{1},E_{2},E_{3}$ each with different purpose. We first show that the digraph induced by $E_{1}$ contains a collection ${\mathcal{C}}_{1}$ of $O(\log n)$ vertex disjoint directed cycles that cover all vertices, i.e. a vertex cycle cover. We then use $E_{2}$ to transform ${\mathcal{C}}_{1}$ into a vertex cycle cover ${\mathcal{C}}_{2}$ in which each cycle $C\in{\mathcal{C}}_{2}$ has size at least $n/\log n$. We use $E_{3}$ to transform ${\mathcal{C}}_{2}$ into a Hamilton cycle.
@@
We use (24) and (25) to obtain (8) from (7). The final inequality uses the fact that $z\sim m/n$ if $m/n\to\infty$. Our bound on $|SMALL|$ follows from the Markov inequality.
(b) The probability that a vertex is part of the randomly chosen edge of an L-switch is at most $\Delta_{0}/m$. So the probability there is some vertex that is chosen twice in this way is at most $n(\Delta_{0}/m)^{2}=o(1)$. Apart from this, a vertex can be involved twice, once in a P-switch and once in an L-switch, w.h.p.
@@
(d) Let $\Delta_{0}=\log^{2}n$ be a high probability bound on the maximum degree, proved in Lemma 12 below. The probability that (d) fails can be bounded by
@@
$\sum_{k=1}^{\ell_{0}}\binom{n}{k+1}k!\binom{k}{3}\binom{m}{k}\left(\frac{\Delta_{0}}{n}\right)^{2k}\left(\left(\frac{3e^{2}\log^{4}n\Delta_{0}}{m}\right)+\sum_{\ell=0}^{\log n/100}\binom{m-k}{\ell}\frac{1}{n^{\ell}}\left(1-\frac{1}{n}\right)^{m-k-\ell}\right)^{3}=o(1).$
@@
We now discuss the construction of $E_{1},E_{2},E_{3}$. Starting with a random ${\bf x}\in\Omega_{1}$ we construct ${\bf y}=y_{1},y_{2},\ldots,y_{2m}$ randomly as in Claim 3. We let $j_{1}=n\log n/5$ and initialise $E_{1}$ as $\left\{(y_{2j-1},y_{2j}):j\leq j_{1}\right\}$. Then we go through y in order $j=n\log n/5+1,\ldots,m$ and add $(y_{2j-1},y_{2j})$ to $E_{1}$ if either $y_{2j-1}$ has appeared less than $\log n/100$ times in $y_{1},y_{3},\ldots,y_{2j_{1}-1}$ or if $y_{2j}$ has appeared less than $\log n/100$ times in $y_{2},y_{4},\ldots,y_{2j_{1}}$. We then randomly partition the edges $E_{\bf y}\setminus E_{1}$ into $E_{2},E_{3}$ and then re-order the edges so that $E_{2}\cap M$ appears before $E_{3}\cap M$ where $M$ is the set of edges induced by $y_{i},i>n\log n/5$.
@@ 2.2 Phase 1
Let $m=\tfrac{n}{2}(\log n+2\log\log n+c_{n})$ satisfy $m/n\lesssim\log n$, then
$\lim_{n\to\infty}\mathbb{P}(G_{1}\text{ has a perfect matching })=\begin{cases}0&c_{n}\to-\infty.\\ e^{-e^{-c}/4}&c_{n}\to c.\\ 1&c_{n}\to\infty.\end{cases}$ (9)
@@
Let ${\mathcal{M}}_{1}$ denote the event that $G_{1}$ contains two vertices of degree one in either $A$ or $B$ that share a common neighbour. Let ${\cal N}_{k},k\geq 3$ denote the event that $G_{1}$ contains a set $S\subseteq A:|S|=k\text{ and }|N_{1}(S)|\leq k-1$. We want a minimal $S$ and so we refine the definition of ${\cal N}_{k}$ by adding (i) $|N_{1}(S)|=k-1$ and (ii) each $b\in N_{1}(S)$ has at least two neighbours in $S$, (we can reduce $S$ by one if there is a $b\in N_{1}(S)$ with one neighbour in $S$. For then we can find a smaller $S$) and (iii) $S\cup N_{1}(S)$ induces a connected subgraph. This reduction must stop before $k=1$ because of our minimum degree condition. Then
@@
Let an A-cherry be a path of length two with centre vertex in $A$ and endpoints of degree one in $B$. We define B-cherries by reversing the roles of A and B. We let $Z$ denote the number of A-cherries. Then in $D_{\bf x}$,
$\mathbb{E}(Z)$ $=n\binom{n}{2}\sum_{d\geq 2}\mathbb{P}(d_{\bf x}^{-}(1)=d_{\bf x}^{-}(2)=1,d_{\bf x}^{+}(3)=d\mid{\mathcal{E}}^{\prime},{\mathcal{E}}^{\prime\prime})\frac{m(m-1)d(d-1)}{m^{2}(m-1)^{2}}$
$\sim\frac{n^{3}}{2m^{2}}\left(\sum_{d=2}^{\Delta_{0}}\frac{z^{d+2}}{(d-2)!(e^{z}-1)^{3}}+\mathbb{P}(d\geq\Delta_{0})\right),$ (11)
removing the conditioning on ${\mathcal{E}}^{\prime},{\mathcal{E}}^{\prime\prime}$ using (24), (25)
$\sim\frac{n^{3}z^{4}}{2m^{2}e^{2z}},\quad\text{ since $\mathbb{P}(d\geq\Delta_{0})=(\log n)^{-\Omega(\log^{2}n)}$ and $e^{z}-1\sim e^{z}$ as $m/n\to\infty$},$
@@
$\geq e^{2\omega}/8\to\infty.$ (14)
@@
At this point we want to use the Chebyshev inequality to prove that $Z\neq 0$ w.h.p. Thus we write $\mathbb{E}(Z(Z-1)=A_{1}+A_{2}$, accounting disjoint A-cherries and non-disjoint A-cherries where $A_{1}$ counts disjoint cherries and $A_{2}$ counts pairs of A-cherries sharing an A-vertex (forming a $K_{1,3}$). In this case, $z\geq(5-o(1))\log n/14$, which is needed for (20) below.
$A_{1}=$ (15)
@@
$\hskip 216.81pt\frac{m(m-1)d_{1}(d_{1}-1)}{m^{2}(m-1)^{2}}\frac{(m-2)(m-3)d_{2}(d_{2}-1)}{(m-2)^{2}(m-3)^{2}}nn$ (16)
$\sim\mathbb{E}(Z)^{2},nn$ (17)
$A_{2}=nn$ (18)
$n\binom{n}{3}\sum_{d\geq 3}\mathbb{P}(d_{\bf x}^{-}(1)=d_{\bf x}^{-}(2)=d_{\bf x}^{-}(3)=1,d_{\bf x}^{+}(4)=d\mid{\mathcal{E}}^{\prime},{\mathcal{E}}^{\prime\prime})\frac{m(m-1)(m-2)d(d-1)(d-2)}{m^{2}(m-1)^{2}(m-2)^{2}}nn$ (19)
@@
Explanation: in the above expression, $d$ is the degree of vertex $b_{1}$ and Lemma 12 below shows that $d\leq\Delta_{0}$ with probability $1-(\log n)^{-\Omega(\log^{2}n)}$. The 2 accounts the common neighbour being in $A$ or $B$. Having chosen $a_{1},a_{2},a_{3},b_{1}$ in at most $n\binom{n}{3}$ ways, we choose the relevant edges in at most $m^{3}$ ways and then $m^{3}d^{3}m^{-6}$ approximates the probability that the edges are as claimed.
@@
$\mathbb{E}((Z)_{k})\sim n^{3k}\sum_{t=0}^{k}\binom{k}{t}m^{2k}n^{-4k}\left(1-\frac{2k}{n}\right)^{m-2k}\sim\left(\frac{e^{-c}}{4}\right)^{k}.$
It follows from the method of moments that $Z$ is asymptotically Poisson with mean $e^{-c}/4$.
@@
Going back to (14), with $c_{n}=\omega$, we see that we see that $\mathbb{E}(Z)\leq e^{-2\omega}/8\to 0$.

### v2

@@ 1 Introduction
$\lim_{n\to\infty}\mathbb{P}(D_{n,m}^{(\delta\geq 1)}\text{ is Hamiltonian})=\begin{cases}0&c_{n}\to-\infty.\\ e^{-e^{-c}/8}&c_{n}\to c.\\ 1&c_{n}\to\infty.\end{cases}$ (2)
@@ 1.1 Previous work
Previous work in this area has focused on $G_{n,m}^{(\delta\geq k)}$, where we condition on $G_{n,m}$ having minimum degree at least $k$. This work was inspired by the result of Komlós and Szemerédi [14]: suppose that $m=\frac{n}{2}(\log n+\log\log n+c_{n})$. Then,
@@ 2 Proof of Theorem 1
We will base our proof on a 3-phase method used in various forms in [6], [8], [10], [12]. In Phase 1 we construct a set of $O(\log n)$ vertex disjoint directed cycles that cover all vertices. In Phase 2, we transform this cover so that each cycle has length at least $n/\log^{1/2}n$. In Phase 3, we patch these cycles together to create a Hamilton cycle. There is also a Phase 0 in which we partition the edge set into $E_{1},E_{2},E_{3}$ for use in the corresponding phase.
We first need to describe a useable model for $D_{n,m}^{(\delta\geq 1)}$. For a sequence ${\bf x}=(x_{1},x_{2},\ldots,x_{2m})\in[n]^{2m}$ we let $D_{\bf x}$ be the multi-digraph with vertex set $[n]$ and edge set
$E_{\bf x}=\left\{(x_{2j-1},x_{2j}):j=1,2,\ldots,m\right\}$. Let $d^{+}_{{\bf x}}(i)=|\left\{j:x_{2j-1}=i\right\}|$ be the out-degree of $i\in[n]$ in $D_{\bf x}$ and similarly let $d^{-}_{{\bf x}}(i)=|\left\{j:x_{2j}=i\right\}|$ be the in-degree of $i\in[n]$ in $D_{\bf x}$. Then $\Omega_{1}=\left\{{\bf x}\in[n]^{2m}:d^{+}_{{\bf x}}(i),d^{-}_{{\bf x}}(i)\geq 1\right\}$ for all $i\in[n]$.
@@
For ${\bf x}\in\Omega_{1}$, let $L_{\bf x}=\left\{j\in[m]:x_{2j-1}=x_{2j}\right\}$ be the set of loops in $D_{\bf x}$, let $M_{\bf x}=\{j\in[m]:\exists j^{\prime}\neq j\text{ s.t. }(x_{2j^{\prime}-1},x_{2j^{\prime}})=$ $(x_{2j-1},x_{2j})\}$ define the set of multiple edges in $D_{\bf x}$; and let $\Delta_{\bf x}=\max_{j}(d_{{\bf x}}^{-}(j)+d_{{\bf x}}^{+}(j))$. (We show later, see Lemma 12 below, that if $m\leq n\log n$ then w.h.p. there are severe restrictions on the closeness and multiplicity of loops and multiple edges.)
@@
Consequently, choosing ${\bf x}$ (near) uniformly from $\Omega_{1}^{*}$ and taking $D_{\bf x}$ is equivalent to sampling (near) uniformly from ${\mathcal{D}}_{n,m}^{(\delta\geq 1)}$. By sampling uniformly from $\Omega_{1}$ and then using switchings to remove any loops or parallel edges we obtain a space $\Omega_{1}^{(0,0)}$ which is almost all of $\Omega_{1}^{*}$ and whose entries are (near) uniformly distributed $\Omega_{1}^{(0,0)}$.
We can generate a random member of $\Omega_{1}$ as follows: Let $Z=Z(z)$ be a truncated Poisson random variable with parameter $z\geq 1$, where
@@
Note that
$z\sim\rho\text{ if }\rho\to\infty.$ (6)
Indeed,
$z\leq\rho=z+\frac{1}{1+\frac{z^{2}}{2!}+\frac{z^{2}}{3!}+\cdots+}\leq z+1.$
Note also that $f(x)=xe^{x}/(e^{x}-1)$ is strictly monotone increasing and $f(0)=1,f(\infty)=\infty$ and so the first equation defines $z$ uniquely. This follows from $(e^{x}-1)^{2}f^{\prime}(x)=e^{2x}-(1+x)e^{x}>0$.
and note
@@
Let $Z_{i}^{\prime},Z_{i}^{\prime\prime},i\in[n]$ be independent copies of a truncated Poisson random variable $Z(z)$. Let ${\bf x}$ be chosen randomly from $\Omega_{1}$. Then $\{d_{\bf x}^{+}(i)\}_{j\in[n]}$ is distributed as $\{Z_{i}^{\prime}\}_{i\in[n]}$ conditional on $\sum_{j\in[n]}Z_{i}^{\prime}=m$ and similarly for $\{d_{\bf x}^{-}(i)\}_{i\in[n]}$ with respect to $Z_{i}^{\prime\prime}$.
@@
Given ${\bf x}\in\Omega_{1}^{(k,0)}$ with $k>0$, we will define an $L$-switch that removes a loop $(x,x)$ by choosing an edge $(a,b)$ and replacing the pair by $(a,x),(x,b)$. More formally, suppose that $i\in L_{\bf x}$. Let $j\notin L_{\bf x}$. Then an $L$-switch produces ${\bf x}^{\prime}$ where we make the replacements $x^{\prime}_{2i-1}\leftarrow x_{2j-1},x^{\prime}_{2j-1}\leftarrow x_{2i-1}$. After $k$ $P$-switches we will have produced a member of $\Omega_{1}^{(0,0)}$. (The above can make new parallel edges, but they are few in number and we show in Lemma 5 that we can exclude switches that create new parallel edges.)
$P$-switches preserve the precise degree sequence $d_{{\bf x}}^{+}(i),d_{{\bf x}}^{-}(i),i\in[n]$ for ${\bf x}\in\Omega_{1}$ but $L$-switches do not. On the other hand, $L$-switches preserve the total degrees $d_{{\bf x}}^{+}(i)+d_{{\bf x}}^{-}(i),i\in[n]$ for ${\bf x}\in\Omega_{1}$. However, they do preserve $d_{{\bf x}}^{+}(i)\geq 1,d_{{\bf x}}^{-}(i)\geq 1$ w.h.p. for $i\in[n]$.
@@
With a little bit of effort, we can use the analysis for Claim 3 to obtain an estimate of the size of ${\mathcal{D}}_{n,m}^{(\delta\geq 1)}$, see Appendix C.
@@ 2.1 Phase 0
Our input will be a sequence y constructed as in Claim 3, starting with a random ${\bf x}\in\Omega_{1}$. We will discuss Hamiltonicity in the context of $D_{\bf y}$. This phase partitions $E_{\bf y}$ into three subsets $E_{1},E_{2},E_{3}$ each with different purpose. We first show that the digraph induced by $E_{1}$ contains a collection ${\mathcal{C}}_{1}$ of $O(\log n)$ vertex disjoint directed cycles that cover all vertices, i.e. a vertex cycle cover. We then use $E_{2}$ to transform ${\mathcal{C}}_{1}$ into a vertex cycle cover ${\mathcal{C}}_{2}$ in which each cycle $C\in{\mathcal{C}}_{2}$ has size at least $n/\log^{1/2}n$. We use $E_{3}$ to transform ${\mathcal{C}}_{2}$ into a Hamilton cycle.
@@
We use Proposition 1 from Appendix B to obtain (9) from (8). The final inequality uses the fact that $z\sim m/n$ if $m/n\to\infty$. Our bound on $|SMALL|$ follows from the Markov inequality.
(b) The probability that a vertex is part of the randomly chosen edge of an $L$-switch is at most $\Delta_{0}/m$. Here, throughout the paper,
$\Delta_{0}=\log^{2}n$
is a high probability bound on the maximum degree, (proved in Lemma 12 below). So the probability there is some vertex that is chosen twice in this way is at most $n(\Delta_{0}/m)^{2}=o(1)$. Apart from this, a vertex can be involved twice, once in a $P$-switch and once in an $L$-switch, w.h.p. See Lemma 12 in the appendix for the constraints on loops and parallel edges.
@@
(d) The probability that (d) fails can be bounded by
@@
$\sum_{k=1}^{\ell_{0}}\binom{n}{k+1}k!\binom{k}{3}\binom{m}{k}\left(\frac{\Delta_{0}}{n}\right)^{2k}\left(\left(\frac{3e^{2}\log^{4}n\Delta_{0}}{m}\right)+\sum_{\ell=0}^{\log n/100}\binom{m-k}{\ell}\frac{1}{n^{\ell}}\left(1-\frac{1}{n}\right)^{m-k-\ell}\right)^{3}\\ =o(1).$
@@
We now discuss the construction of $E_{1},E_{2},E_{3}$. Starting with a random ${\bf x}\in\Omega_{1}$ we construct ${\bf y}=y_{1},y_{2},\ldots,y_{2m}$ randomly as in Claim 3. We let $j_{1}=n\log n/5$ and initialise $E_{1}$ as $\left\{(y_{2j-1},y_{2j}):j\leq j_{1}\right\}$. Then we go through y in order $j=n\log n/5+1,\ldots,m$ and add $(y_{2j-1},y_{2j})$ to $E_{1}$ if either $y_{2j-1}$ has appeared less than $\log n/100$ times in $y_{1},y_{3},\ldots,y_{2j_{1}-1}$ or if $y_{2j}$ has appeared less than $\log n/100$ times in $y_{2},y_{4},\ldots,y_{2j_{1}}$. We randomly place the edges $E_{\bf y}\setminus E_{1}$ into $E_{2},E_{3}$, each with probability 1/2, and then re-order the edges so that $E_{2}\cap M$ appears before $E_{3}\cap M$ where $M$ is the set of edges induced by $y_{i},i>n\log n/5$. This re-ordering is not very important, it just helps in our description of the constructions. We just need two large sets of sufficiently random edges.
@@ 2.2 Phase 1
Let $m=\tfrac{n}{2}(\log n+2\log\log n+c_{n})$ satisfy $m/n\lesssim\log n$, where $A_{n}\lesssim B_{n}$ if $A_{n}\leq(1+o(1)B_{n}$ as $n\to\infty$. (This bounds $c_{n}$ and for larger $c_{n}$ we have minimum in- and out-degree at least one w.h.p.) Then
$\lim_{n\to\infty}\mathbb{P}(G_{1}\text{ has a perfect matching })=\begin{cases}0&c_{n}\to-\infty.\\ e^{-e^{-c}/8}&c_{n}\to c.\\ 1&c_{n}\to\infty.\end{cases}$ (10)
@@
Let ${\mathcal{M}}_{1}$ denote the event that $G_{1}$ contains two vertices of degree one in either $A$ or $B$ that share a common neighbour. Let ${\cal N}_{k},k\geq 3$ denote the event that $G_{1}$ contains a set $S\subseteq A:|S|=k$ such that $|N_{1}(S)|\leq k-1$. A minimal such $S$ satisfies (i) $|N_{1}(S)|=k-1$ and (ii) each $b\in N_{1}(S)$ has at least two neighbours in $S$, (we can reduce $S$ by one if $|N_{1}(S)|<k-1$ or if there is a $b\in N_{1}(S)$ with one neighbour in $S$) and (iii) $S\cup N_{1}(S)$ induces a connected subgraph. If ${\mathcal{M}}_{1}$ does not occur then this reduction in the size of $S$ must stop before $k=1$ because of our minimum degree condition. Then, because $G_{1}$ has no isolated vertices,
@@
Let an A-cherry be a path of length two with centre vertex in $A$ and endpoints of degree one in $B$. We define B-cherries by reversing the roles of A and B. We let $Z$ denote the number of A-cherries. Then in $D_{\bf x}$, with ${\mathcal{E}}^{\prime},{\mathcal{E}}^{\prime\prime}$ as in (7), ${\mathcal{D}}_{d}=\left\{d_{\bf x}^{-}(1)=d_{\bf x}^{-}(2)=1,d_{\bf x}^{+}(3)=d\right\}$,
$\mathbb{E}(Z)$ $=n\binom{n}{2}\sum_{{\mathcal{D}}_{d}\geq 2}\mathbb{P}({\mathcal{D}}_{d}\mid{\mathcal{E}}^{\prime},{\mathcal{E}}^{\prime\prime})\frac{m(m-1)d(d-1)}{m^{2}(m-1)^{2}}$
$=n\binom{n}{2}\sum_{d\geq 2}\frac{\mathbb{P}\left({\mathcal{D}}_{d},\sum_{i\geq 3}Z_{i}^{\prime\prime}=m-2,\sum_{i\geq 2}Z_{i}^{\prime}=m-d\right)}{\mathbb{P}\left(\sum_{i}Z_{i}^{\prime\prime}=m,\sum_{i}Z_{i}^{\prime}=m\right)}\frac{d(d-1)}{m(m-1)}$
$=n\binom{n}{2}\sum_{d\geq 2}\frac{\mathbb{P}({\mathcal{D}}_{d})\mathbb{P}\left(\sum_{i\geq 3}Z_{i}^{\prime\prime}=m-2\right)\mathbb{P}\left(\sum_{i\geq 2}Z_{i}^{\prime}=m-d\right)}{\mathbb{P}\left(\sum_{i}Z_{i}^{\prime\prime}=m\right)\mathbb{P}\left(\sum_{i}Z_{i}^{\prime}=m\right)}\frac{d(d-1)}{m(m-1)}$
$\sim n\binom{n}{2}\sum_{d\geq 2}\mathbb{P}({\mathcal{D}}_{d})\cdot\frac{d(d-1)}{m(m-1)}$ (12)
after removing the probabilities of sums of $Z_{i}^{\prime},Z_{i}^{\prime\prime}$ using Proposition 1 in Appendix B
$\sim\frac{n^{3}}{2m^{2}}\left(\sum_{d=2}^{\Delta_{0}}\frac{z^{d+2}}{(d-2)!(e^{z}-1)^{3}}+\mathbb{P}(d_{\bf x}^{+}(3)\geq\Delta_{0})\right),$ (13)
$\sim\frac{n^{3}z^{4}}{2m^{2}e^{2z}},\quad\text{ since $\mathbb{P}(d_{\bf x}^{+}(3)\geq\Delta_{0})=(\log n)^{-\Omega(\log^{2}n)}$ and $e^{z}-1\sim e^{z}$ as $m/n\to\infty$},$
@@
$\geq e^{\omega}/8\to\infty.$ (15)
It is worth observing that the above shows that conditioning on ${\mathcal{E}}^{\prime},{\mathcal{E}}^{\prime\prime}$ only changes probabilities by a factor $1+o(1)$ and from now on we rely on this without mention.
@@
At this point we want to use the Chebyshev inequality to prove that $Z\neq 0$ w.h.p. Thus we write $\mathbb{E}(Z(Z-1))=A_{1}+A_{2}$, where $A_{1}$ counts disjoint cherries and $A_{2}$ counts pairs of A-cherries sharing an A-vertex (forming a $K_{1,3}$). In this case, $z\geq(5-o(1))\log n/14$, which is needed for (16) below.
$A_{1}=$
@@
$\hskip 216.81pt\frac{m(m-1)d_{1}(d_{1}-1)}{m^{2}(m-1)^{2}}\frac{(m-2)(m-3)d_{2}(d_{2}-1)}{(m-2)^{2}(m-3)^{2}}$
$)^{2},\qquad\text{ using \eqref{EZ}}.$ ( ) .
$A_{2}=$
$n\binom{n}{3}\sum_{d\geq 3}\mathbb{P}(d_{\bf x}^{-}(1)=d_{\bf x}^{-}(2)=d_{\bf x}^{-}(3)=1,d_{\bf x}^{+}(4)=d\mid{\mathcal{E}}^{\prime},{\mathcal{E}}^{\prime\prime})\frac{m(m-1)(m-2)d(d-1)(d-2)}{m^{2}(m-1)^{2}(m-2)^{2}}$
@@
Explanation: in the above expression, $d$ is the degree of vertex $b_{1}$ and Lemma 12 below shows that $d\leq\Delta_{0}$ with probability $1-(\log n)^{-\Omega(\log^{2}n)}$. The 2 accounts for the common neighbour being in $A$ or $B$. Having chosen $a_{1},a_{2},a_{3},b_{1}$ in at most $n\binom{n}{3}$ ways, we choose the relevant edges in at most $m^{3}$ ways and then $m^{3}d^{3}m^{-6}$ approximates the probability that the edges are as claimed.
@@
$\mathbb{E}((Z)_{k})\sim n^{3k}\sum_{t=0}^{k}\binom{k}{t}m^{2k}n^{-4k}\left(1-\frac{2k}{n}\right)^{m-2k}\sim\left(\frac{e^{-c}}{8}\right)^{k}.$
It follows from the method of moments that $Z$ is asymptotically Poisson with mean $e^{-c}/8$.
@@
Going back to (15), with $c_{n}=\omega$, we see that we see that $\mathbb{E}(Z)\leq e^{-\omega}/8\to 0$.